Using the 5-step method to solve:
A sailboat travels 20 miles downstream in 4 hours. It returns in 5 hours. Find the speed of the sailboat in still water and the rate of the current.

Answer by mananth(16946)

boat speed = x
current speed = y
..
distance = 20 miles
time = d/r
..
downstream speed = 20/(x+y)
upstream speed = 20/(x-y)
..
20/(x+y)=4
4x+4y=20
/4
x+y=5...............1
20/(x-y)=5
5x-5y=20
/5
x-y=4...............2
1 x + 1 y = 5 .............1
1 x + -1 y = 4 .............2
Add the two equations
2 x = 9
/ 2
x = 4.50
plug value of x in (1)
1 x + 1 y = 5
4.5 + 1 y = 5
1 y = 5 -4.5
1 y = 0.5
y = 0.50